# Kuaiji: the First Chinese Accounting Large Language Model

### 21 Feb 2024 v1

@@ Abstract
Large Language Models (LLMs) like ChatGPT and GPT-4 have demonstrated impressive proficiency in comprehending and generating natural language. However, they encounter difficulties when tasked with adapting to specialized domains such as accounting. To address this challenge, we introduce Kuaiji, a tailored Accounting Large Language Model. Kuaiji is meticulously fine-tuned using the Baichuan framework, which encompasses continuous pre-training and supervised fine-tuning processes. Supported by CAtAcctQA, a dataset containing 70,000 genuine accountant-client dialogues, Kuaiji exhibits exceptional accuracy and response speed. Our contributions encompass the creation of the first Chinese accounting dataset, the establishment of Kuaiji as a leading open-source Chinese accounting LLM, and the validation of its efficacy through real-world accounting scenarios.
@@ 1 Introduction
  • To the best of our knowledge, we have created the inaugural Chinese accounting dataset, featuring 70,000 real instances spanning 14 distinct accounting subareas.
  • Kuaiji stands as the premier open-source Chinese accounting LLM designed explicitly for accounting tasks, serves as the huge knowledge base and inference tools.
@@ 2 Related Work
Large Language Models. The advent of Large Language Models (LLMs), exemplified by breakthroughs like ChatGPT OpenAI (2022) and GPT-4 OpenAI (2023), has ignited a transformative wave in artificial intelligence research. These models have not only revolutionized natural language understanding and generation but have also paved the way for groundbreaking advancements across diverse applications. Despite the proprietary nature of some LLMs’ training methodologies, the emergence of open-source alternatives such as Baichuan Yang et al. (2023a), LLaMA Touvron et al. (2023), Bloom Workshop et al. (2023), and Falcon Penedo et al. (2023) has democratized access to state-of-the-art language modeling capabilities. Generative LLMs, or Language Models, exhibit remarkable capabilities in generating text that is both coherent and contextually relevant. Through pretraining on extensive text corpora and fine-tuning for alignment with human instructions, they excel in producing human-like text in response to given prompts or inputs. In essence, LLMs operate by modeling the probability of a sentence $s$ (comprising a sequence of word tokens $q_{1},q_{2},\ldots,q_{n}$) as $p(s)=\prod_{i}^{n}p(q_{i}\mid q_{<i})$. Here, $q_{i}$ represents the $i$-th token of the sentence $s$, while $q_{<i}$ denotes the partial word token sequence preceding the $i$-th step.
@@ 3 Dataset Construction
In this section, we present an exposition on the datasets utilized for training the Kuaiji model. To enhance the efficacy of model training, we meticulously curated two distinct datasets tailored for various stages of training: a continuous pre-training dataset and a supervised fine-tuning dataset. These datasets encompass fundamental and specialized knowledge spanning disciplines such as accounting, auditing, financial management, economics, and economic law. Rigorous manual curation and validation were employed to ensure the quality of these datasets, resulting in the procurement of a high-quality corpus.
@@ 3.1 Construction of Continuous Pre-training Dataset
Diversification of the dataset was achieved through comprehensive data collection efforts spanning various channels. These channels included specialized financial investigation reports, annual reports of publicly listed companies, national statistical bureaus, financial commentaries from columns, media coverage of financial news, and other open datasets. The diverse array of data sources not only enhances the dataset’s breadth but also imbues it with real-time relevance, mirroring the complexities of real-world scenarios. Moreover, these sources serve as complementary resources to traditional academic materials, fostering a comprehensive understanding of professional knowledge within the model.
In tandem with efforts to enhance dataset diversity, rigorous data cleansing procedures were implemented to ensure the dataset’s quality and integrity. This involved meticulous error correction, noise reduction, and consistency checks. Additionally, stringent measures were taken to safeguard data privacy, including the removal of sensitive information and anonymization techniques. Furthermore, expert validation from professionals spanning accounting, economics, law, and related fields was sought to provide an additional layer of scrutiny, thereby upholding the dataset’s professionalism and accuracy. The types and sizes of the datasets we finally obtained are presented in Table 1.
@@ 3.2 Construction of Supervised Fine-Tuning Dataset
Fine-tuning is crucial for the training of large models, as it enables models to better understand and generate text relevant to specific domains or tasks by training on domain-specific data. However, due to the scarcity of fine-tuning datasets in certain domains, we opted to utilize domain-specific practice exercises and exam papers to aid in dataset construction. Nevertheless, the existing pool of practice exercises may not cover all necessary aspects comprehensively. To address this shortfall, we devised a specialized prompt to assist in generating the fine-tuning dataset with the help of GPT.
@@ 3.2.1 From Test
Similarly, expert validation was conducted to ensure the clarity and conciseness of the questions, minimizing ambiguity and ensuring the explicit understandability of the tested content. We prepared a total of 15,677 data points derived from these practice question collections.
3.2.2 From GPT
We selected our self-constructed Continuous Pre-training dataset and supplemented it with fine-tuning data generated by GPT. Initially, we designed a prompt (as depicted in Table 2) to automatically generate instructions, input, and output by analyzing the input text and formatting the output accordingly. Additionally, we implemented a sliding window mechanism to ensure that GPT reads the text with a fixed window size (in this case, the number of lines) and slides at an appropriate distance. This approach guarantees that the generated content is contextually relevant, thereby enhancing logical coherence, consistency, and accuracy while minimizing the risk of semantic discontinuity. Through comparative testing, we determined that the optimal window size and sliding stride are 50 and 30, respectively.
Table 2: Instruction Construction Prompt.
@@ 4 Methodology
This section delves into the fine-tuning process of Kuaiji, spanning two key stages: Continuous Pre-Training (CPT) in Section 4.1 and Supervised Fine-Tuning (SFT)in Section 4.2. Each stage is discussed sequentially to mirror the research workflow. Additionally, we provide a comprehensive method flowchart, depicted in Figure 1, to offer a visual representation of the methodology employed in Kuaiji.
@@ 4.2 Supervised Fine-Tuning
Supervised fine-tuning is a pivotal step in refining the Kuaiji model for specific accounting tasks. Leveraging task-specific labeled datasets from both human and strong LLMs, such as the meticulously curated CAtAcctQA dataset and the distilled data from GPT, this phase focuses on adapting the pre-trained model to the intricacies of accounting language and domain-specific conventions. With high-quality domain dialogue data, the model can effectively invoke the accounting knowledge accumulated during pre-training, thereby understanding and responding to users’ queries. SFT involves adjusting the model’s parameters through gradient-based optimization techniques, guided by the labeled data’s ground truth annotations. By iteratively feeding batches of labeled examples to the model and updating its weights through back-propagation, the model learns to refine its internal representations to better align with the target accounting tasks. This process allows Kuaiji to capture nuances in accounting terminology, understand contextual cues unique to financial documents, and effectively handle domain-specific queries and tasks. The supervised fine-tuning phase is instrumental in enhancing Kuaiji’s accuracy, responsiveness, and overall performance in accounting language processing, making it well-suited for real-world applications in finance and accounting domains.
Moreover, to avoid the collapse of inherent capabilities rather than learning substantive ones Gudibande et al. (2023); Shumailov et al. (2023), we mix the domain-specific instructions with general ones, with the proportion of 80%:20% Chen et al. (2023); Dong et al. (2024).
@@ 4.3 Training Strategy: QLoRA
The 4-bit NormalFloat quantization technique introduced in QLoRA Dettmers et al. (2023) is based on Quantile Quantization and Block-wise Quantization. It involves mapping the parameters of the neural network to a 4-bit representation while preserving the distribution of the original parameters. This is achieved by dividing the parameter space into quantiles and mapping each quantile to a specific value.
Using the components described above, QLoRA for a single linear layer in the quantized base model with a single LoRA adapter Hu et al. (2021) is defined as follows:
$Y^{\text{BF16}}=X^{\text{BF16}}\cdot\text{doubleDequant}(c^{\text{FP32}},c^{k\text{-bit}},W^{\text{NF4}})+$ (1)
@@
Where $\text{doubleDequant}(\cdot)$ is defined as:
@@
$(c^{\text{FP32}},c^{k\text{-bit}}),W^{\text{4bit}}\bigl{)}=W^{\text{BF16}}.$
5 Experiments
@@ 5.1 Training Details
Our model is based on Baichuan2-13B-chat, a specialized Chinese Large Language Model (LLM) consisting of 13 billion parameters, specifically designed for conversational tasks. The training process utilizes 4 A100-80G GPUs with parallelization techniques within the Baichuan2 framework. During training, we implement the low-rank adaptation (LORA) parameter-efficient tuning method with rank equal to 8 to optimize model performance. This methodology is facilitated through the Transformers4 and PEFT5 libraries. To ensure efficient resource utilization, we adopt fp16 precision with ZeRO-2, along with gradient accumulation strategies. Additionally, to manage computational complexity, we limit the length of individual responses to 4096 tokens. Optimization techniques include the AdamW optimizer, a dropout rate of 0.1, and a cosine learning rate scheduler. A portion of the training dataset (10%) is reserved for validation, and the best-performing checkpoints are selected as the final model. Mechanisms are also in place to handle loss during gradient explosions and to adjust the learning rate for training stability. The hyperparameters for each stage of training are iteratively refined to ensure convergence within an effective loss range.

### 24 Feb 2024 v2

@@ Abstract
Large Language Models (LLMs) like ChatGPT and GPT-4 have demonstrated impressive proficiency in comprehending and generating natural language. However, they encounter difficulties when tasked with adapting to specialized domains such as accounting. To address this challenge, we introduce Kuaiji, a tailored Accounting Large Language Model. Kuaiji is meticulously fine-tuned using the Baichuan framework, which encompasses continuous pre-training and supervised fine-tuning processes. Supported by CAtAcctQA, a dataset containing 15,677 genuine accountant-client dialogues, Kuaiji exhibits exceptional accuracy and response speed. Our contributions encompass the creation of the first Chinese accounting dataset, the establishment of Kuaiji as a leading open-source Chinese accounting LLM, and the validation of its efficacy through real-world accounting scenarios.
@@ 1 Introduction
  • To the best of our knowledge, we have created the inaugural Chinese accounting dataset, featuring 15,677 real instances spanning 14 distinct accounting subareas.
  • Kuaiji stands as the premier open-source Chinese accounting LLM designed explicitly for accounting tasks and serves as a huge knowledge base and inference tools.
@@ 2 Related Work
Large Language Models. The advent of Large Language Models (LLMs), exemplified by breakthroughs like ChatGPT OpenAI (2022) and GPT-4 OpenAI (2023), has ignited a transformative wave in artificial intelligence research. These models have not only revolutionized natural language understanding and generation but have also paved the way for groundbreaking advancements across diverse applications. Despite the proprietary nature of some LLMs’ training methodologies, the emergence of open-source alternatives such as Baichuan Yang et al. (2023a), LLaMA Touvron et al. (2023), Bloom Scao et al. (2022); Workshop et al. (2023), and Falcon Penedo et al. (2023) has democratized access to state-of-the-art language modeling capabilities. Generative LLMs, or Language Models, exhibit remarkable capabilities in generating text that is both coherent and contextually relevant. Through pretraining on extensive text corpora and fine-tuning for alignment with human instructions, they excel in producing human-like text in response to given prompts or inputs. In essence, LLMs operate by modeling the probability of a sentence $s$ (comprising a sequence of word tokens $q_{1},q_{2},\ldots,q_{n}$) as $p(s)=\prod_{i}^{n}p(q_{i}\mid q_{<i})$. Here, $q_{i}$ represents the $i$-th token of the sentence $s$, while $q_{<i}$ denotes the partial word token sequence preceding the $i$-th step.
@@ 3 Dataset Construction
In this section, we will introduce the datasets used for training the Kuaiji model. To enhance the model’s training effectiveness, we constructed a Continuous Pre-training (CPT) dataset and a Supervised Fine-Tuning (SFT) dataset tailored for different training stages. For the construction of the CPT dataset, we carefully considered both the professionalism and breadth of the accounting domain. Through an in-depth search for professional-oriented data and extensive acquisition of related domain data, we curated a CPT dataset totaling 828MB in size (cf. Section  3.1). Similarly, in preparing the SFT dataset, we reorganized existing data to meet our dataset requirements and leveraged the assistance of existing strong Language Models (LLMs) to handle labor-intensive divergent data preparation tasks. This process culminated in a dataset structured into instructions, inputs, and outputs, containing 35,784 fine-tuning data entries (cf. Section  3.2).
We’ve organized our dataset following the criteria of the Chinese Certified Public Accountants (CPA) into six main categories: Accounting, Auditing, Corporate Strategy and Risk Management, Tax Laws, Economic Laws, and Financial Management and Cost Management. Additionally, there’s a miscellaneous category labeled Others in Figure 1.
Figure 1: Statistics on the distribution of Kuaiji training dataset.
[…]
@@ 3.1 Construction of Continuous Pre-training Dataset
Diversification of the dataset was achieved through comprehensive data collection efforts spanning various channels. These channels included specialized financial investigation reports [^2] https://www.ctrchina.cn/, annual reports of publicly listed companies [^3] https://www.cninfo.com.cn/, national statistical bureaus  [^4] https://www.stats.gov.cn/, financial commentaries from columns [^5] https://finance.eastmoney.com/, media coverage of financial news  [^6] https://finance.sina.com.cn/, and other open datasets  [^7] https://www.chinairn.com/. The diverse array of data sources not only enhances the dataset’s breadth but also imbues it with real-time relevance, mirroring the complexities of real-world scenarios. Moreover, these sources serve as complementary resources to traditional academic materials, fostering a comprehensive understanding of professional knowledge within the model.
In tandem with efforts to enhance dataset diversity, rigorous data cleansing procedures were implemented to ensure the dataset’s quality and integrity. This involved meticulous error correction, noise reduction, and consistency checks. Additionally, stringent measures were taken to safeguard data privacy, including the removal of sensitive information and anonymization techniques. Furthermore, expert validation from professionals spanning accounting, economics, law, and related fields was sought to provide an additional layer of scrutiny, thereby upholding the dataset’s professionalism and accuracy.
In the end, we divided the obtained data into two categories: specialized and related domains, and mixed them in a ratio of 7:3. This mixture helps to expand the training dataset, accelerate the convergence speed of the model, and enhance its robustness. The types and sizes of the datasets are presented in Table 1.
@@ 3.2 Construction of Supervised Fine-Tuning Dataset
Fine-tuning is crucial for the training of large models, as it enables models to better understand and generate text relevant to specific domains or tasks by training on domain-specific data. Due to the scarcity of fine-tuning datasets in certain domains, we opted to utilize domain-specific practice exercises and exam papers to aid in dataset construction (cf. Section 3.2.1). However, because exam questions mainly focus on concept and scenario questions, there is a lack of understanding in more fields. Although the quality of the dataset constructed through exam questions is excellent, collecting exam questions is extremely labor-intensive and the amount of data is limited. To address this shortfall, we devised a specialized prompt following Self-Instruct Prompt Wang et al. (2022) to assist in generating the fine-tuning dataset (cf. Section 3.2.2) with the help of strong LLMs given reference domain text, i.e. GPT. The constructed instructions can be referred to in Table 3.
@@ 3.2.1 From Test
Similarly, expert validation was conducted to ensure the clarity and conciseness of the questions, minimizing ambiguity and ensuring the explicit understandability of the tested content. We prepared a total of 3,497 instructions derived from these practice question collections.
3.2.2 From Strong LLMs
We selected our self-constructed Continuous Pre-training dataset and supplemented it with fine-tuning data generated by Strong LLMs, i.e. GPT, as they are trained under huge and various general knowledge, such as high-quality literature corpus and dialogues, etc., which is beneficial for comprehending the accounting knowledge and generating new dialogues without domain-specific knowledge.
Initially, we crafted the self-instruction prompt, as illustrated in Table 2, to automate the generation of instructions, inputs, and outputs. This process involves analyzing the input domain text and formatting the output accordingly, using the domain-specific text dataset as a reference. To facilitate efficient text analysis, we introduced a sliding window mechanism. This mechanism allows GPT to comprehend the domain text within a fixed window size—measured in lines—and slide over the text by a predetermined distance to continue reading. It is important to highlight that we directed the LLMs to strictly follow given Requirements and Reference Formats, aiming to produce instructions that are not only diverse but also detailed.
This methodology ensures the relevance of the generated content to the context, significantly improving logical coherence, consistency, and accuracy, while also reducing the likelihood of semantic discontinuities. Through comparative analyses, we determined the ideal window size and sliding stride to be 50 lines and 30 lines, respectively. This optimization enhances the effectiveness of our approach in generating contextually relevant and coherent content.
Table 2: Self-Instruction Prompt.
@@ 3.2.2 From Strong LLMs
Table 3: Fine-tuning Data Example.
[Fine-tuning Data Example]
{ “Instruction”: “List the ways in which investment properties can be converted and the corresponding accounting treatment, and briefly explain the principles of conversion.”,
Input”: “Method of conversion of investment properties, Accounting treatment, Explanation of conversion principles”,
Output”: “An enterprise that converts investment property means that the enterprise has conclusive evidence of a change in the use of its real estate, and should convert investment property to other assets if one of the following conditions is met, for example, if the investment property begins to be used for its own purposes. In the cost model, the book value of the real estate before conversion should be used as the recorded value after conversion. When an investment property measured by the fair value model is converted to real estate for its own use, the fair value on the date of conversion shall be used as the book value of the real estate for its own use. The difference between the fair value and the original book value shall be recognized in profit or loss for the current period. Under the cost model, in case of conversion of investment property to owner-occupied property, the corresponding accounting treatment is:
Debit: Fixed assets (intangible assets)
Accumulated depreciation (accumulated amortization) of investment properties
Provision for impairment of investment properties
Credit: Investment properties - cost
Accumulated depreciation (accumulated amortization)
Provision for impairment of fixed assets (intangible assets)
The conversion principles are as follows:
1. Comparability principle: The conversion form and conversion date should be consistent with previous accounting periods to ensure the comparability of financial statements.
2. Reliability principle: The conversion form and date should be based on reliable evidence and data to ensure the accuracy and reliability of the conversion process.
3. Consistency principle: Once a certain conversion form and conversion date are selected, it should be consistent in future accounting periods unless there are sufficient reasons for change.
4. Dominance principle: The selection of the conversion form and conversion date should be based on the dominant characteristics and economic substance of investment real estate to reflect its true value and financial condition.
5. Economic practicability principle: The selection of the conversion form and conversion date should take into account economic practicability to meet users’ information needs and decision-making purposes.” }
@@ 4 Methodology
This section delves into the fine-tuning process of Kuaiji, spanning two key stages: Continuous Pre-Training (CPT) in Section 4.1 and Supervised Fine-Tuning (SFT) in Section 4.2. Each stage is discussed sequentially to mirror the research workflow. Additionally, we provide a comprehensive method flowchart, depicted in Figure 3, to offer a visual representation of the methodology employed in Kuaiji. The whole training framework adopts QLoRA strategy as stated in Section 4.3.
@@ 4.2 Supervised Fine-Tuning
Supervised fine-tuning is a pivotal step in refining the Kuaiji model for specific accounting tasks. Leveraging task-specific labeled datasets from both human and strong LLMs, such as the meticulously curated CAtAcctQA dataset and the distilled data from GPT, this phase focuses on adapting the pre-trained model to the intricacies of accounting language and domain-specific conventions. With high-quality domain dialogue data, the model can effectively invoke the accounting knowledge accumulated during pre-training, thereby understanding and responding to users’ queries. SFT involves adjusting the model’s parameters through gradient-based optimization techniques, guided by the labeled data’s ground truth annotations. By iteratively feeding batches of labeled examples to the model and updating its weights through back-propagation, the model learns to refine its internal representations to better align with the target accounting tasks. This process allows Kuaiji to capture nuances in accounting terminology, understand contextual cues unique to financial documents, and effectively handle domain-specific queries and tasks. The supervised fine-tuning phase is instrumental in enhancing Kuaiji’s accuracy, responsiveness, and overall performance in accounting language processing, making it well-suited for real-world applications in finance and accounting domains. Moreover, to avoid the collapse of inherent capabilities rather than learning substantive ones Gudibande et al. (2023); Shumailov et al. (2023), we mix the domain-specific instructions with general ones, with the proportion of 80%:20% Chen et al. (2023); Dong et al. (2024).
We have conducted statistics on the token count and logarithmic frequency of the queries and answers of the instructions we generated, and the results are shown in Figure 2.
@@ 4.3 Training Strategy: QLoRA
In the modified version of QLoRA that incorporates 8-bit NormalFloat quantization, based on Quantile Quantization and Block-wise Quantization, the parameters of the neural network are mapped to an 8-bit representation. This enhances the precision by offering 256 possible values for each parameter, compared to the 16 values provided by the original 4-bit representation. The main goal remains to preserve the distribution of the original parameters by dividing the parameter space into quantiles, each mapped to a specific 8-bit value. The definition of QLoRA for a single linear layer in the quantized base model with a single LoRA adapter, adjusted for 8-bit quantization, is as follows:
$Y^{\text{BF16}}=X^{\text{BF16}}\cdot\text{doubleDequant}(c^{\text{FP32}},c^{k\text{-bit}},W^{\text{NF8}})+$ (1)
@@
Where $\text{doubleDequant}(\cdot)$ is adapted for 8-bit quantization as:
@@
$(c^{\text{FP32}},c^{k\text{-bit}}),W^{\text{8bit}}\bigl{)}=W^{\text{BF16}}.$
In adopting 8-bit quantization (NF8) for weight parameters $W$, we aim to strike an optimal balance between enhanced precision and efficient memory use, with the possibility of adjusting $W$ and $c_{2}$ block sizes tailored to specific application needs and computational constraints. Priority is given to updating adapter weight gradients $\frac{\partial E}{\partial L_{i}}$, thus sidestepping the computation of gradients for the 8-bit quantized weights $\frac{\partial E}{\partial W}$. The process involves de-quantizing weights from their NF8 storage representation to the BF16 computational format to compute the derivative $\frac{\partial X}{\partial W}$ with BFloat16 precision, ensuring precision in gradient calculation without excessive memory demand.
5 Empirical Analysis
@@ 5.1 Training Details
Our model is based on Baichuan2-13B-chat, a specialized Chinese Large Language Model (LLM) consisting of 13 billion parameters, specifically designed for conversational tasks. The training process utilizes 4 A100-80G GPUs with parallelization techniques within the Baichuan2 framework. During training, we implement the low-rank adaptation (LORA) parameter-efficient tuning method with a rank equal to 8 to optimize model performance. This methodology is facilitated through the Transformers4 and PEFT5 libraries. To ensure efficient resource utilization, we adopt fp16 precision with ZeRO-2, along with gradient accumulation strategies. Additionally, to manage computational complexity, we limit the length of individual responses to 4096 tokens. Optimization techniques include the AdamW optimizer, a dropout rate of 0.1, and a cosine learning rate scheduler. A portion of the training dataset (10%) is reserved for validation, and the best-performing checkpoints are selected as the final model. Mechanisms are also in place to handle loss during gradient explosions and to adjust the learning rate for training stability. The hyperparameters for each stage of training are iteratively refined to ensure convergence within an effective loss range.
During the model training process, we set the learning rate to 5e-05 and utilized a cosine annealing scheduler to adjust the learning rate dynamically. Regarding gradient clipping, we applied a threshold of 1.0 to prevent gradient explosion and enhance training stability. To control resource wastage during training, we constrained the maximum input length to 512 tokens. This approach is feasible because the majority of information is typically contained within the initial few tokens of a sequence. Limiting the maximum length ensures that the model focuses more on crucial information, thereby mitigating overfitting and unnecessary computations.
@@ 5.2 Baselines
5.3 Case Study
To demonstrate the effectiveness of accounting, we compared Kuaiji and two other LLMs in three case studies.
Table 4: Case study 1.
Query
How are the three accounting statements linked in the valuation model?
ChatGPT
The income statement, balance sheet, and cash flow statement are interconnected in the valuation model through adjustments and transactions that flow between them, such as net income affecting retained earnings on the balance sheet and operating cash flows derived from net income on the cash flow statement.
[…]
The three accounting statements linked in the valuation model are the income statement, the balance sheet, and the cash flow statement.
[…]
In the valuation model, there is a close relationship between the three accounting statements (i.e. balance sheet, income statement, and cash flow statement). Firstly, the balance sheet reflects the assets, liabilities, and owner’s equity of a company, reflecting its financial condition and debt paying ability. The income statement reflects the company’s revenue and expenses, including its profit and loss situation. The cash flow statement reflects the company’s cash inflows and outflows, including cash flows from investments, financing, and operating activities. The connection between these three statements is that they all reflect the financial condition and operating results of the company, but use different financial indicators and accounting methods. By integrating these three reports together, investors can have a comprehensive understanding of the company’s financial condition and profitability, enabling more accurate valuation analysis.
Table 5: Case study 2.
Query
Assuming a tax rate of 25%, how do the three statements change when one hundred dollars of new depreciation is added?
ChatGPT
Adding \$100 of new depreciation would decrease taxable income by \$100, resulting in a tax savings of \$25 (25% of \$100).
[…]
Adding \$100 of new depreciation will result in a reduction in pre-tax income, a higher depreciation asset on the balance sheet, and a decrease in operating cash flow, offset by an increase in investing cash flow.
[…]
When a new depreciation of \$100 is added, all three reports will undergo changes. Firstly, in the balance sheet, the book value of fixed assets will decrease by \$100, and accumulated depreciation will increase by \$100. Secondly, in the income statement, depreciation expenses will increase by \$100, and the impact on profit depends on whether depreciation expenses belong to expensed depreciation. If the depreciation expense belongs to expensed depreciation, the total profit will increase by \$100; If depreciation expenses belong to capitalized depreciation, the total profit will not change significantly. Finally, in the cash flow statement, depreciation expenses will increase by \$100, affecting the cash flows generated from operating activities. Therefore, an increase of \$100 in new depreciation will affect the company’s financial condition, operating results, and cash flow, and the impact on the financial statements needs to be comprehensively considered.
Table 6: Case study 3.
Query
If you could only use two accounting statements to analyze a company, which two would you choose? Why?
ChatGPT
I would choose the income statement and the cash flow statement because they provide essential insights into a company’s profitability and cash flow dynamics, offering a comprehensive view of its financial performance and sustainability.
[…]
If I could only use two accounting statements to analyze a company, I would choose the income statement and the balance sheet. Both statements provide valuable information about the company’s financial health and performance.
[…]
If you can only use two financial statements to analyze a company, I would choose the balance sheet and cash flow statement. The reasons are as follows: 1. The balance sheet can reflect the financial condition of a company, including assets, liabilities, and owner’s equity, which helps us understand the company’s asset structure and financial strength. Through the balance sheet, we can calculate financial indicators such as the company’s debt-to-equity ratio and asset turnover to evaluate the company’s debt-paying ability and operational efficiency. The cash flow statement reflects the company’s cash inflows and outflows, including cash flows from investments, financing, and operating activities. Through the cash flow statement, we can understand the company’s cash sources and purposes, as well as the cash flow situation. This helps us evaluate the company’s cash liquidity and cash flow from operating activities. Therefore, the balance sheet and cash flow statement can provide us with the financial and cash flow situation of the company, and the combination of these two statements can help us comprehensively evaluate the financial and operating conditions of the company.
Figure 4: Kuaiji sigle-turn Q&A test (in Chinese).
Figure 5: Kuaiji multi-turns Q&A test (in Chinese).
5.4 Application Scenarios
After training on a large number of accounting professional datasets, our accounting model greatly assists in the following subfields:
  1. 1.
    Financial Analysis and Reporting: Kuaiji can analyze company financial statements, provide interpretations of financial indicators, and identify potential risks and opportunities. This is crucial for businesses to understand their financial health, make informed decisions, and communicate effectively with stakeholders.
  2. 2.
    Tax Consultancy: Kuaiji can offer guidance and advice on tax laws, assisting businesses in compliance and optimizing tax strategies. Given the complexity of tax regulations, having a reliable tool like Kuaiji can ensure accurate tax filings and minimize the risk of penalties or audits.
  3. 3.
    Risk Management: Utilizing Kuaiji, companies can identify and assess financial risks, receive risk management recommendations, and reduce their exposure to risks. By proactively managing risks, businesses can safeguard their assets, reputation, and long-term viability.
  4. 4.
    Management Decision Support: Kuaiji can analyze and interpret financial data, provide decision support to management, and aiding in strategic planning, and business implementation. With Kuaiji’s insights, executives can make data-driven decisions, allocate resources effectively, and drive organizational growth.
  5. 5.
    Audit and Compliance: Kuaiji can assist in auditing processes, identifying anomalies and potential non-compliance issues, and ensuring adherence to relevant regulations and standards. This helps companies maintain transparency, integrity, and trust among investors, regulators, and the public.
  6. 6.
    Education and Training: Kuaiji can serve as an educational tool, offering learning resources and guidance for students in accounting and finance, aiding in the understanding of complex accounting principles and practices. By leveraging Kuaiji, educators can enhance the learning experience and equip future professionals with practical skills.
  7. 7.
    Market Forecasting and Investment Decision Making: Kuaiji can analyze financial data and market trends to forecast future market movements, providing guidance for investment decisions. In a volatile market environment, having accurate predictions from Kuaiji can help investors mitigate risks and capitalize on opportunities.
  8. 8.
    Financial Planning and Consultancy: Kuaiji can provide customized financial planning and advice based on individual or business financial situations, helping them achieve their financial goals. Whether it’s retirement planning, debt management, or wealth accumulation, Kuaiji can offer personalized strategies for financial success.
Our Kuaiji can achieve multiple rounds of dialogue and retain memory function. We chose an accounting calculation question to test Kuaiji’s response in Figure 4, and in addition, we also chose a multi-round dialogue task to demonstrate Kuaiji’s effectiveness in Figure 5.